Calculate the number of pounds of 

CO2 released into the atmosphere when a 20.0 gallon

20.0 gallon tank of gasoline is burned in an automobile engine. Assume that gasoline is primarily octane, C


8

H


18

,

C8H18, and that the density of gasoline is 0.692 g⋅mL

−1

.

0.692 g⋅mL−1. This assumption ignores additives. Also, assume complete combustion.


1
Expert's answer
2021-10-20T01:40:02-0400

Density of the gasoline ,d= 0.692 g/mL


Volume of gasoline in an tanks,V = 20.0 gallons = 75708.2 mL


Let mass of the gasolin be MM = V × d = 75708.2 mL × 0.692 g/mL=52390.0744g

"2C_8H_{18}+25O_2\rightarrow 16CO_2+18H_2O"

Mass of octane burnt in the tank = M = 57,629.081 g

Moles of octane ="\frac{52390.0744g}{114.08 g/mol}=459.24mol"

According to reaction, 2 moles of octane gives 16 moles of carbon-dioxide.


Then 505.1637 mol of octane will give

"\frac{16}{2}\times 459.24mol=3673.92mol of carbon-dioxide"

Mass of 3673.92 mol of carbon-dioxide:

3673.92mol × 44.01 g/mol = 161689.22 g


Mass of carbon-dioxide produced in pounds =

0.00220462× 161689.22= 356.46pounds

356.46pounds of carbon-dioxide released into the atmosphere